OpenStudy (anonymous):

A video store charges a monthly membership fee of $7.50, but the charge to rent each movie is only $1.00 per movie. Another store has no membership fee, but it costs $2.50 to rent each movie. The equation below represents this situation where m is the number of movies rented each month. 7.50 + 1.00m = 2.50m Which of the following is the number of movies that need to be rented for the total fees to be the same from either store? A.3 Movies B.5 Movies C.7 Movies D.9 Movies

OpenStudy (anonymous):

just plug in the numbers, and then solve. such as: 7.50+1.00(3)= 2.50(3) 7.50+3.00=7.50 11.50=7.50 does that make sense??

OpenStudy (anonymous):

No, i got confused. Did yu multiply by 3 ?

OpenStudy (anonymous):

yeah wherever the letter is, plug in one of hte numbers from the selection given...

OpenStudy (anonymous):

Do you get it now??

OpenStudy (anonymous):

Yes I did Is B = 5

OpenStudy (anonymous):

yup. that looks right..